OpenStudy (anonymous):

Larry drew a blue marble from a bag and, without replacing it, drew a red marble from the same bag. Are these events dependent or independent? A. dependent B. independent

OpenStudy (anonymous):

Define the word dependent in your textbook

OpenStudy (anonymous):

look, does the probability change? in other words, does P(Red on 1st draw) = P(Red on 2nd draw|1st was Blue)? if so, then the events are independent, if not, they're dependent.

OpenStudy (anonymous):

The dependent variable is the variable calculated from the independent variable.

OpenStudy (anonymous):

no different def.

OpenStudy (anonymous):

requiring someone or something for financial or other support.

OpenStudy (anonymous):

suppose there are 3 red and 3 blue in the bag. what is the probability of drawing a red marble?

OpenStudy (anonymous):

i think 3 or 18

OpenStudy (anonymous):

?

OpenStudy (anonymous):

sorry but i am bad at math

OpenStudy (anonymous):

i think you're just goofing off

OpenStudy (anonymous):

keep telling yourself that and you will continue to be

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the probability of any event is always between 0 and 1, inclusive

OpenStudy (anonymous):

k

OpenStudy (anonymous):

is it depented

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

what if the marbles are replaced?

OpenStudy (anonymous):

if 2 events, A & B, are independent then \(P(A\cap B) = P(A)\cdot P(B)\)